Can anyone give me the solution to this in y+mx+b format?

8x-7y-14=0


Thank you

Answer by fractalier(2101)

Sure...we solve this for y...so we get
8x - 7y - 14 = 0
8x - 7y = 14
-7y = -8x + 14
y = (8/7)x - 2